Table 1. Average Return and Beta
Stock 1 2 3 4 5
Average annual return 15% 8% 12% 50% 80%
Beta 1 1.2 2.5 3 3.5
Table 2. Portfolio A information
Stock 1 2 3
Investment (million USD)

If you manage a portfolio worth $100,000,000 and you choose to invest in only stock 1, 2, and 3.

What is the minimum beta for a portfolio with return equal to 13%?

No short sales allowed.

Report the answer with 2 numbers after decimal place, say 1.51

**Please show excel formulas**
Answer #1

To get the portfolio return of 13% there will be the following weightage of stocks is as follows

Stock 1 = 45%

Stock 2 = 8%

Stock 3 = 47%

Portfolio Return = 45*0.15 + 8*0.08+47*0.12 = 13 %

So

Portfolio beta  = 1*45%+1.2*8%+2.5*47% = 1.721